OpenStudy (driftracer305):

express as single log :- 4 log x - 6 log (x+2)

OpenStudy (driftracer305):

@ganeshie8

OpenStudy (aum):

\(\Large A\log(B) = \log(B)^A\) \(\Large \log(P) - \log(Q) = \log(\frac PQ) \)

OpenStudy (anonymous):

\[\log\]\[x^4\]/\[(x+2)^2\]

OpenStudy (anonymous):